# Search for Quadruplet Scalars using Boosted Decision Trees at the LHC

### v1

@@ 4 Analysis and Results
with S and B being the number of signal and background events, respectively, after the imposition of the BDT cut. Essentially, it quantifies the signal’s distinguishability from expected background noise. The significance of the signal S1 and S2 are listed in Table 6. We observe that the discovery prospect (i.e., $\mathcal{S}\geq 5$) for the scalars through the process S1 lies up to approximately 600 to 700 GeV, while for exclusion $\mathcal{S}\gtrsim 2$ level, the limit can be extended up to $\sim 900$ GeV. We have found that the significance is larger in S2 compared to S1, even though the cross-section is smaller in S2. Using S2, we can exclude ($\mathcal{S}\gtrsim 2$) a new scalar up to a mass of approximately 1 TeV (1000 GeV) or more.
@@
We would also like to mention that the lepton multiplicity distribution ($N_{lep}$), as shown in Figure 5 indicates that a requirement $N_{lep}\geq 5$ effectively suppresses the SM background processes. We have checked that a similar analysis carried in this channel suppress B1, B2, and B3 to negligible levels. The major contribution in the background comes from B4, which has cross-section similar to $S1$ at 500 GeV mass. as the primary remaining contamination. Therefore, we find that the cut significantly reduce background contamination, leading to a noticeable improvement in signal significance in both S1 and S2. We find that with this revised search strategy, the discovery prospect for S1 reaches up to $\sim 600-700$ GeV. while the exclusion limit reaches $\sim 900$ GeV. However, the signal cross-section is also significantly reduced compared to the previous analysis. With so few events, reconstruction of either the fermion or the scalar states is not feasible. Therefore, we do not pursue this channel in further detail.
@@ 5 Summary
We explore the prospective collider signatures in these two scenarios, alongside the prevailing experimental constraints from existing collider searches. We found that the signals with 4 or more leptons along with 2 jets offer substantial signal $\times$ BR and are not excluded by the LHC experiment. We study the pair production mode $\phi_{1}^{+}\phi_{1}^{-}$ and associated production mode $\phi^{++}\phi_{2}^{-}$ and their decays via the fermion quintuplets at $14$ TeV LHC, giving rise to final states with 4 or more leptons along with 2 jets. We focus on the fermiophilic scenario and demonstrate that scalar and fermion masses can be reconstructed through a selection strategy based on kinematic variables, applied to the final-state particles.
We also conduct a detailed collider analysis for the final state with at least four leptons and two jets, incorporating SM background events within a multivariate analysis (MVA) framework, to assess the discovery prospects of these scalars. The pair production process for the scalars achieves a sensitivity $5\sigma$ with a mass of the scalar ranging from $\approx 600$ to $700\text{ GeV}$ and an exclusion limit ($2\sigma$) up to $\approx 900\text{ GeV}$. Critically, the associated production process offers superior sensitivity, extending the exclusion potential ($2\sigma$) for the scalar mass to $1\text{ TeV}$ and beyond. The signal processes feature cascade decays that pose significant challenges to reconstruction. We have also demonstrated that the successful reconstruction of quadruplet scalars and quintuplet fermions is feasible in both channels, S1 and S2.
We also examine the discovery and exclusion reach in the $5\ell+2$ jets final state. Although this channel yields even higher significance, the number of surviving signal events after all selections falls below 10 at $\mathcal{L}=3000$ fb-1. Furthermore, for scalar masses exceeding 1 TeV, the signal cross-section decreases substantially, rendering discovery unattainable in these final states. Also, at higher energies, the $W$ bosons become highly energetic and could be reconstructed as fat jets; however, such an analysis lies beyond the scope of this work.

### v2

@@ 4 Analysis and Results
with S and B being the number of signal and background events, respectively, after the imposition of the BDT cut. Essentially, it quantifies the signal’s distinguishability from expected background noise. The significance of the signal S1 and S2 are listed in Table 6. We observe that the discovery prospect (i.e., $\mathcal{S}\geq 5$) for the scalars through the process S1 lies up to approximately 600 to 700 GeV, while for exclusion $\mathcal{S}\gtrsim 2$ level, the limit can be extended up to $\sim 900$ GeV. We have found that the significance is larger in S2 compared to S1, even though the cross section is smaller in S2. Using S2, we can exclude ($\mathcal{S}\gtrsim 2$) a new scalar up to a mass of approximately 1 TeV (1000 GeV) or more.
@@
We would also like to mention that the lepton multiplicity distribution ($N_{lep}$), as shown in Figure 5 indicates that a requirement $N_{lep}\geq 5$ effectively suppresses the SM background processes. We have checked that a similar analysis carried in this channel suppress B1, B2, and B3 to negligible levels. The major contribution in the background comes from B4, which has cross section similar to $S1$ at 500 GeV mass. as the primary remaining contamination. Therefore, we find that the cut significantly reduce background contamination, leading to a noticeable improvement in signal significance in both S1 and S2. We find that with this revised search strategy, the discovery prospect for S1 reaches up to $\sim 600-700$ GeV. while the exclusion limit reaches $\sim 900$ GeV. However, the signal cross section is also significantly reduced compared to the previous analysis. With so few events, reconstruction of either the fermion or the scalar states is not feasible. Therefore, we do not pursue this channel in further detail.
@@ 5 Summary
We explore the prospective collider signatures in these two scenarios, alongside the prevailing experimental constraints from existing collider searches. We find that the signals with 4 or more leptons along with 2 jets offer substantial signal $\times$ BR and are not excluded by the LHC experiment. We study the pair production mode $\phi_{1}^{+}\phi_{1}^{-}$ and associated production mode $\phi^{++}\phi_{2}^{-}$ and their decays via the fermion quintuplets at $14$ TeV LHC, giving rise to final states with 4 or more leptons along with 2 jets. We focus on the fermiophilic scenario and demonstrate that scalar and fermion masses can be reconstructed through a selection strategy based on kinematic variables, applied to the final-state particles.
We also conduct a detailed collider analysis for the final state with at least four leptons and two jets, incorporating SM background events within a multivariate analysis (MVA) framework, to assess the discovery prospects of these scalars. The pair production process for the scalars achieves a sensitivity of $5\sigma$ with mass of the scalar ranging from $\approx 600$ to $700\text{ GeV}$ and an exclusion limit ($2\sigma$) up to $\approx 900\text{ GeV}$. Critically, the associated production process offers superior sensitivity, extending the exclusion potential ($2\sigma$) for the scalar mass to $1\text{TeV}$ and beyond. The signal processes feature cascade decays that pose significant challenges to reconstruction. We have also demonstrated that the successful reconstruction of quadruplet scalars and quintuplet fermions is feasible in both channels, S1 and S2.
We also examine the discovery and exclusion reach in the $5\ell+2$ jets final state. Although this channel yields even higher significance, the number of surviving signal events after all selections falls below 10 at $\mathcal{L}=3000$ fb-1. Furthermore, for scalar masses exceeding 1 TeV, the signal cross section decreases substantially, rendering discovery unattainable in these final states. Also, at higher energies, the $W$ bosons become highly energetic and could be reconstructed as fat jets; however, such an analysis lies beyond the scope of this work.